OpenStudy (anonymous):

7 is what percent of 3500?

OpenStudy (anonymous):

woody causes 9/11

OpenStudy (homeworksucks):

so, 7parts/3500, is how many parts of 100

OpenStudy (homeworksucks):

so you can set up a ratio

OpenStudy (homeworksucks):

\[\frac{ 7 }{ 3500 } = \frac{ x }{ 100 }\]

OpenStudy (homeworksucks):

Have you done basic algebra yet?

OpenStudy (homeworksucks):

Cross multiply, which gives you 700=3500x, and divide

OpenStudy (anonymous):

so 5?

OpenStudy (homeworksucks):

Yea

OpenStudy (anonymous):

wait isnt it .05%?

OpenStudy (homeworksucks):

Neither, my bad, its 0.2%, 700/3500=0.2

OpenStudy (anonymous):

it okay thanks